An investment with CF₀= - $120,000 and CFᴊ= +20,000 (J= 1,2,………,10). If the interest rate is 4% compounded annually,

  1. Compute the NPV of the investment
  2. Is the investment economically acceptable ? why ?
  3. Find the ROR.
Answer #1

NPV is 42218 , yes the investment is economically acceptable as the NPV is positive and Return on investment is more than the interest rate.

ROR of Investment is 16.667%

Calculation and Explanation

Net present Value ( NPV ) is the present value of future cashflow at a given discount rate .

Formula for NPV = -Initial Cash outflow + Present value of Cash flows

CF0= Intial cash flow at point 0 =$12000

CFj = Anual Cash Inflow (as Positive ) where J is 1 to 10 year =20000 for 10 year

Discount Rate =r =4% compunding annually

NPV= -120000+ $20000×PVAF (4%,10)

NPV =-$120000+$20000×8.110895731

NPV =-1200000+162217.9146

NPV =$42217.91 or $42218

PVAF = Present value annuity factor for 4%, for 10 year , ( refer present value annuity table if provided or u can calculate it in excel by compount interest Pv formula 1/(1+r)n and then adding all 10 year value

( if any issue which this calculation please ask on comment section )

Rate of Return ( ROR ) is the return on the Average capital employed

Net Income ÷ capital employed or Intial Investment

As Nothing is provided about the Net income or non cash expenses it assume that the cashflow is the only income

ROR =20000/120000= 16.667 %